OpenStudy (anonymous):

The dimensions of four different cylinders are given below. Which cylinder has the greatest volume? Cylinder 1:radius equals 4 feet, height equals 3 feet Cylinder 2:radius equals 2 feet, height equals 5 feet Cylinder 3:radius equals 3 feet, height equals 4 feet Cylinder 4:radius equals 5 feet, height equals 1 feet Cylinder 1 Cylinder 2 Cylinder 3 Cylinder 4

jhonyy9 (jhonyy9):

do you know formula for volum of a cylinder ?

OpenStudy (anonymous):

use the formula of volume for cylinder and calculate yourslef: \[V = \frac{ 1 }{ 3 }\pi \times r^{2} \times h\]